OpenStudy (anonymous):

ok, here's another on lol. (sqrt5 + 1)(7sqrt5 + 1). i got 15sqrt5 + 1 as my answer, is this correct?

OpenStudy (saifoo.khan):

\[36+8\sqrt{5}\]

OpenStudy (saifoo.khan):

that's the answer!!

OpenStudy (saifoo.khan):

Did you used FOIL?

OpenStudy (anonymous):

i thought i did lol

OpenStudy (anonymous):

when i did FOIL, i got 7sqrt5 + sqrt5 + 7sqrt5 + 1

OpenStudy (saifoo.khan):

\[\Large 7\sqrt{5} \times \sqrt{5} = 35\]

OpenStudy (anonymous):

i didn't multiply within the sqrt.......... o.O *smacking head*

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

now i have 36 + 8sqrt5 lol. i have one last problem lol

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):